How do I find the ordered pairs for this equation 2x+y=10 I am not sure how to start I think I SOLVE X AND THEN Y.aM I CORRECT?
Answer by funmath(2933)

These are kind of cool because you get to pick the points to solve for. You get to make this as easy as you can.
:
2x+y=10
Let x=0 and solve for y: (0 is easy!)
2(0)+y=10
0+y=10
y=10 This ordered pair (x,y)=(0,10)
:
Now let's let y=0 and solve for x:
2x+0=10
2x=10
2x/2=10/2
x=5 This ordered pair (x,y)=(5,0)
:
Let x=1 and solve for y: (1 is easy too!)
2(1)+y=10
2+y=10
-2+2+y=10-2
0+y=8
y=8 This ordered pair is (1,8)
:
What easy numbers do you want x to be? There are an infinite number of points to choose, make it as easy as you can on yourself.
Happy Calculating!!!